OpenStudy (anonymous):

Marvin currently has a balance of $1,192 in an account he has held for 14 years. He opened the account with an initial deposit of $800. What is the simple interest on the account? 2.3% 3.5% 7.1% 10.6%

OpenStudy (anonymous):

Amt - Principal = SI = Principal x Rate x Time / 100

OpenStudy (anonymous):

S.I = amount - principal = 1192 - 800 = 392 = (prt)/100

OpenStudy (anonymous):

It's 3.5 % ?